Records at Hal’s Accounting Services show the following costs for year 1.

Direct materials and supplies $ 50,000
Employee costs 2,900,000
Total overhead 1,340,000

Production was 35,000 billable hours. Fixed overhead was $790,000.

Assuming no change in billable hours in year 2, direct materials and supplies costs are expected to increase by 10 percent. Direct labor costs are expected to increase by 15 percent. Variable overhead per billable hour is expected to remain the same, but fixed overhead is expected to increase by 5 percent.


Required:

a. Year 2 production is expected to be 28,000 billable hours. What are the estimated direct materials, direct labor, variable overhead, and fixed overhead costs for year 2?

b. Determine the total costs per billable hour for year 1 and year 2.
Answer #1

a.

Direct material = [($50,000 / 35,000) + 10%] X 28,000 = $44,000

Direct labor = [($2,900,000 / 35,000) + 15%] X 28,000 = $2,668,000

Variable overhead = [($1,340,000 - $790,000) / 35,000] X 28,000 = $440,000

Fixed overhead = $790,000 + 5% = $829,500

Total cost = $3,981,500

b.

Year 1 = $4,290,000 / 35,000 = $122.5714285

Year 2 = $3,981,500 / 28,000 = $142.1964285